OpenStudy (anonymous):

integral x/(x-2)dx i know u=x-2, making du=dx i tried rewriting as u+2/u but got nowhere.

OpenStudy (anonymous):

Hmm, would this work? \[f(x) = x*u^{-1}\]

OpenStudy (bahrom7893):

u can't do that Moss's brother..andrew u were tryin the right approach, i think u made a mistake somewhere: (u+2)/u = u/u + 2/u = 1 + 2/u Now if u integrate that, u + 2Ln|u|, plug in u=x-2 x-2+2 Ln|x-2| <= Your answer..

OpenStudy (bahrom7893):

Or x - 2 +Ln(x-2)^2+C (no need for absolute value as ^2 is always positive, and I forgot the C earlier) Please click on become a fan if I helped! Thanks!